1. Assume today is December 31, 2016. Barrington Industries expects that its 2017 after-tax operating income [EBIT(1 – T)] will be $430 million and its 2017 depreciation expense will be $70 million. Barrington's 2017 gross capital expenditures are expected to be $100 million and the change in its net operating working capital for 2017 will be $30 million. The firm's free cash flow is expected to grow at a constant rate of 5% annually. Assume that its free cash flow occurs at the end of each year. The firm's weighted average cost of capital is 8.6%; the market value of the company's debt is $2.25 billion; and the company has 180 million shares of common stock outstanding. The firm has no preferred stock on its balance sheet and has no plans to use it for future capital budgeting projects. Using the free cash flow valuation model, what should be the company's stock price today (December 31, 2016)? Round your answer to the nearest cent. Do not round intermediate calculations.

Ans: $_________ per share
Answer #1

Expected EBIT*(1-t) = 430

Depreciation = 70 M

Change in NWC = 30 M

Capex = 100

Annual Cash flows = EBIT*(1-t) + Depreciation - Capex - Change in NWC

= 430 + 70 - 30 - 100 = 370 M

Growth rate in cash flows , g = 5%

WACC , k = 8.6%

Value of the firm = FCF*(1+g)/(k-g) = 370*(1+5%)/(8.6% - 5%) = 10,791.67= 10.79 B

Value of Equity = Firm value - Value of debt = 10.79 - 2.25 = $ 8.54 B

Number of stock outstanding = 180 Million

Stock price today = 8.54 B/ 180 M = $ 47.45 per share